# LSEAttention is All You Need for Time Series Forecasting

### v6

@@ Abstract
Transformer-based architectures have achieved remarkable success in natural language processing and computer vision. However, their performance in multivariate long-term forecasting often falls short compared to simpler linear baselines. Previous research has identified the traditional attention mechanism as a key factor limiting their effectiveness in this domain.To bridge this gap, we introduce LATST, a novel approach designed to mitigate entropy collapse and training instability—common challenges in Transformer-based time series forecasting. We rigorously evaluate LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models. Notably, LATST manage to achieve competitive performance with fewer parameters than some linear models on certain datasets, highlighting its efficiency and effectiveness.
@@ 3.2 Baseline
Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from those used in PatchTST, with a batch size fixed at 8 for simplicity. Rest of the hyparameters could be referred in Tab. 2 where it demonsteates clear pattern between the set of parameters and the categories.
Table 1: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons. The best results are highlighted in bold. LATST denotes the LSEAttention Time-Series Transformer model utilized in this study, while most results for other models are sourced from Ilbert et al. [7] with some ECL result from Liu et al. [16].
Methods LATST SAMformer PatchTST Transformer TSMixer iTrans Dlinear
ECL 96 0.130 0.155 - 0.182 0.173 - 0.140
192 0.151 0.168 - 0.202 0.204 - 0.153
336 0.169 0.183 - 0.212 0.217 - 0.169
720 0.197 0.219 - 0.238 0.242 - 0.203
Traffic 96 0.369 0.407 0.462 0.420 0.409 0.395 0.410
192 0.386 0.415 0.466 0.441 0.637 0.417 0.423
336 0.417 0.421 0.482 0.501 0.747 0.433 0.436
720 0.455 0.456 0.514 0.468 0.688 0.467 0.466
Weather 96 0.146 0.197 0.177 0.227 0.214 0.174 0.176
192 0.196 0.235 0.225 0.256 0.231 0.221 0.220
336 0.249 0.276 0.278 0.278 0.279 0.278 0.265
720 0.317 0.334 0.354 0.353 0.343 0.358 0.323
ETTh1 96 0.381 0.381 0.414 0.509 0.398 0.386 0.375
192 0.411 0.409 0.460 0.535 0.426 0.441 0.405
336 0.433 0.423 0.501 0.570 0.435 0.487 0.439
720 0.426 0.427 0.500 0.601 0.498 0.503 0.472
ETTh2 96 0.280 0.295 0.302 0.396 0.308 0.297 0.289
192 0.343 0.340 0.388 0.413 0.352 0.380 0.383
336 0.380 0.350 0.426 0.414 0.360 0.428 0.448
720 0.399 0.391 0.431 0.424 0.409 0.427 0.605
ETTm1 96 0.299 0.329 0.329 0.384 0.336 0.334 0.299
192 0.342 0.353 0.367 0.400 0.362 0.377 0.335
336 0.375 0.382 0.399 0.461 0.391 0.426 0.369
720 0.431 0.429 0.454 0.463 0.450 0.491 0.425
ETTm2 96 0.174 0.181 0.175 0.200 0.211 0.180 0.167
192 0.230 0.233 0.241 0.273 0.252 0.250 0.224
336 0.280 0.285 0.305 0.310 0.303 0.311 0.281
720 0.376 0.375 0.402 0.426 0.390 0.412 0.397
Exchange 96 0.102 0.161 0.088 0.292 0.343 0.086 0.081
192 0.211 0.246 0.176 0.372 0.342 0.177 0.157
336 0.408 0.368 0.301 0.494 0.484 0.331 0.305
720 0.865 1.003 0.901 1.323 1.204 0.847 0.603
Table 2: Classifications of Dataset where Low Variates: N = 4, D = 32, F = 64; Mid Variates: N = 8 D = 32, F = 128; High Variates: N = 16, D = 128, F = 256; Ultra-High Variates: N = 16 D = 256, F = 512. It demonstrates a clear pattern between number of variates and number of heads, as well as the dimension sizes.
Dataset Type
@@ 3.3 Performance Analysis
Table 3: Performance of LATST with different activation functions for prediction length of 96.
@@
ETTh2 0.290 0.292
ETTm1 0.295 0.305
ETTm2 0.176 0.191
Exchange Rate 0.098 0.115
3.4 Ablation Study
@@
Table 4: Performance of LATST without GeLU compared to Transformer for prediction length of 96.
Dataset Without GeLU Transformer
ETTh2 0.312 0.396
ETTm1 0.333 0.384
ETTm2 0.183 0.200
Exchange Rate 0.105 0.292

### v8

@@ Abstract
Transformer-based architectures have achieved remarkable success in natural language processing and computer vision. However, their performance in multivariate long-term forecasting often falls short compared to simpler linear baselines. Previous research has identified the traditional attention mechanism as a key factor limiting their effectiveness in this domain. To bridge this gap, we introduce LATST, a novel approach designed to mitigate entropy collapse and training instability—common challenges in Transformer-based time series forecasting. We rigorously evaluated LATST across multiple real-world multivariate time series datasets, demonstrating its ability to outperform existing state-of-the-art Transformer models. Notably, LATST, a single-layer Transformer, manages to achieve competitive performance with fewer parameters than some linear models on average, highlighting its efficiency and effectiveness.
@@ 3.2 Baseline
Hyperparameters: For all experiments, the hyperparameter settings for LATST are derived from SAMFomer, with a batch size fixed at 8 for simplicity. Rest of the hyparameters could be referred in Tab. 1 where it demonstrates clear pattern between the set of parameters and the categories.
Table 1: Classifications of Datasets with corresponding combination of hyperparameters where all are single-layer with Low Variates: $N$ = 4, $D$ = 32, $F$ = 64; Mid Variates:$N$ = 8 $D$ = 32, $F$ = 128; High Variates: $N$ = 16, $D$ = 128, $F$ = 256; Ultra-High Variates: $N$ = 16 $D$ = 256, $F$ = 512. It demonstrates a clear positive relationship between number of variates and number of heads, as well as the dimension sizes to better capture the channel interaction.
Dataset Type Single-Layer, LR: 0.0001
@@
Table 2: Performance metric - MSE (mean squared error) for various time series forecasting models across multiple datasets and prediction horizons. The best results are highlighted in bold. LATST denotes the LSEAttention Time-Series Transformer model utilized in this study, while most results for other models are sourced from Ilbert et al. [7] with some ECL results from triplet layers Transformer on average from PatchTST et al. [6] and iTransformer et al. [16].
Methods LATST SAMformer PatchTST Transformer TSMixer iTransformer Dlinear
ECL 96 0.129 0.155 0.129 0.182 0.173 0.148 0.140
192 0.150 0.168 0.147 0.202 0.204 0.162 0.153
336 0.168 0.183 0.163 0.212 0.217 0.178 0.169
720 0.205 0.219 0.197 0.238 0.242 0.225 0.203
Traffic 96 0.369 0.407 0.462 0.420 0.409 0.395 0.410
192 0.386 0.415 0.466 0.441 0.637 0.417 0.423
336 0.397 0.421 0.482 0.501 0.747 0.433 0.436
720 0.433 0.456 0.514 0.468 0.688 0.467 0.466
Weather 96 0.146 0.197 0.177 0.227 0.214 0.174 0.176
192 0.193 0.235 0.225 0.256 0.231 0.221 0.220
336 0.246 0.276 0.278 0.278 0.279 0.278 0.265
720 0.314 0.334 0.354 0.353 0.343 0.358 0.323
ETTh1 96 0.381 0.381 0.414 0.509 0.398 0.386 0.375
192 0.415 0.409 0.460 0.535 0.426 0.441 0.405
336 0.434 0.423 0.501 0.570 0.435 0.487 0.439
720 0.471 0.427 0.500 0.601 0.498 0.503 0.472
ETTh2 96 0.280 0.295 0.302 0.396 0.308 0.297 0.289
192 0.346 0.340 0.388 0.413 0.352 0.380 0.383
336 0.384 0.350 0.426 0.414 0.360 0.428 0.448
720 0.423 0.391 0.431 0.424 0.409 0.427 0.605
ETTm1 96 0.299 0.329 0.329 0.384 0.336 0.334 0.299
192 0.341 0.353 0.367 0.400 0.362 0.377 0.335
336 0.367 0.382 0.399 0.461 0.391 0.426 0.369
720 0.420 0.429 0.454 0.463 0.450 0.491 0.425
ETTm2 96 0.174 0.181 0.175 0.200 0.211 0.180 0.167
192 0.238 0.233 0.241 0.273 0.252 0.250 0.224
336 0.287 0.285 0.305 0.310 0.303 0.311 0.281
720 0.369 0.375 0.402 0.426 0.390 0.412 0.397
Exchange 96 0.102 0.161 0.088 0.292 0.343 0.086 0.081
192 0.201 0.246 0.176 0.372 0.342 0.177 0.157
336 0.395 0.368 0.301 0.494 0.484 0.331 0.305
720 0.935 1.003 0.901 1.323 1.204 0.847 0.603
@@ 3.3 Performance Analysis
Table 3: Performance of LATST with different activation functions for prediction length of 720.
@@
ECL 0.205 0.205
Traffic 0.433 0.437
Weather 0.314 0.314
ETTh1 0.471 0.471
ETTh2 0.423 0.426
ETTm1 0.420 0.423
ETTm2 0.369 0.369
Exchange Rate 0.935 1.043
3.4 Ablation Study Note Updated Yet
@@
Table 4: Performance of LATST without GeLU compared to Transformer for prediction length of 192.
Dataset With GeLU Without GeLU
ECL 0.150 0.153
Traffic 0.386 0.393
Weather 0.193 0.204
ETTh1 0.415 0.433
ETTh2 0.346 0.378
ETTm1 0.341 0.357
ETTm2 0.238 0.246
Exchange Rate 0.201 0.236
Table 5: Overall performance improvement and average size of parameters with data of other models sourced from Ilbert et al. [7]
LATST SAMformer
MSE Error Percentage 107.17% 100.00%
Average Size of Parameter 271.4K 173.2K